OpenStudy (anonymous):

David has a collection of computer games. Last year, he had 400 CDs in his collection and this year he has 125% as many CDs as he had last year. How many CDs does he have now?

OpenStudy (anonymous):

whats the answer i have a bunch of questions

OpenStudy (anonymous):

do you want the answers, or to know how to do them?

OpenStudy (anonymous):

1.25*400=500

OpenStudy (anonymous):

Last month, Jenny’s math test score was 75. This month, her math test score is 93.What percent of last month’s score is this month’s score? A. 124% whats the answer

OpenStudy (anonymous):

sorry about the a 124% dont worry about it

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

75/93 * 100 =80.65% =81%

OpenStudy (anonymous):

no cause the suggestions are a 124% b 135% c 140% d 168%?

OpenStudy (anonymous):

Last month, Jenny’s math test score was 75. This month, her math test score is 93.What percent of last month’s score is this month’s score? whats the answer

OpenStudy (anonymous):

oh sori had mis interpreted the question 0.75x=0.93 x=0.93/0.75 x=1.24 1.24*100=124%

OpenStudy (anonymous):

Casey's grandmother sent her an $80 bookstore gift card for her birthday last year. This year, she sent her a $110 bookstore gift card. Which proportion can be used to find the percent of last year's gift that this year's gift is, if n represents the unknown percent? whats the answer

OpenStudy (anonymous):

0.8x=1.1 x=1.1/0.8 x=1.375 1.375*100 =137.5%